A shell is fired vertically upwards with a velocity v1 from trolley moving horizontally with velocity v2 .a person on the ground observes the motion of the shell as a parabola , whose horizontal range is

Grade:12th pass

2 Answers

Arun
6 years ago
Horizontal range equals:
x = V2 *t
where t is of motion.
Time of motion can be found from (object moving if his height >0) :
ℎ = V1*t −
1/2* g* t^2 = 0
Therefore, time of motion equals:
t = 2*V1/g
Finally, horizontal range equals:
x = V2*t = 2 V1* V2/g
Abhay
5 years ago
Horizontal range =v^2sin2theta/g 
So,
We can write the above equations as ,
Range =2(vsin theta)(vcos theta)/g
Here ,as you can see (vsin theta )=vertical velocity
And........                       (vcos theta)=horizontal velocity.
And ,vertical velocity =v1
          Horizontal velocity =v2
So put these velocities in the horizontal range equation 
The final answer is 2v1v2/g.